### v3

@@ 3 System Architecture
Our approach comprises two main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through SC specifications. The explainability component provides natural language explanations by implementing RAG, thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
@@
Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability. Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above-mentioned components.
@@ 3.1 Black Box Recorder
Our first component focuses on recording accountable information that captures essential messages generated during the navigation task of an autonomous agent. The selected data will guarantee the reproducibility of the mentioned task, including details like the robot’s poses, the employed map, odometry data, linear and angular velocity, and laser scan readings.
@@
The immutability of the previously described information is achieved through blockchain. Blockchain technology usually faces limitations regarding the size of the data that can be written into a single transaction. Efficiently storing autonomous agent messages becomes challenging due to these constraints. Our proposed solution involves selectively storing the hash of messages at intervals, providing in addition the chance of using different storage frequencies depending on the topic and its relevance. This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
@@
Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the SC specification. This section delves into their design and justifies how both align with the constraints of autonomous agents.
@@
Key features such as consistency are guaranteed by the blockchain’s consensus protocol, where the content of each block is hashed and stored in the next one. Additionally, availability is also assured through the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
3.1.1 Chain of Integrity Proofs
@@
To preserve the integrity of the messages without compromising system performance, our solution entails building a chain where each message hash is linked with the previous hash value. This approach ensures that any alteration of the selected messages will result in a different final hash, easing tampering detection. The verification process is also simplified since only the last hash in the chain needs to be recomputed and compared with the last stored value. A single modification of a hash will disrupt the entire chain from that point onward, simplifying the identification of tampering.
@@
We define a chain entry ($CE$) with a counter $i\in 1..m$ to be linked with the previous one by concatenating the prior digest $h_{i-1}$ with the current log message $Msg_{i}$. hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, commonly used for data integrity verification and digital signatures.
@@
The initial hash value $h_{0}$ of the chain is calculated by applying the selected hash function to a set of $l=32$ random bytes ($RB$), to generate a 256-bit value. Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that integrity proofs for similar initial log messages from different records are not identical. Data privacy is further prioritized by storing in the blockchain only the hashes, rather than the contents of the messages.
@@
By chaining every hash with the previous one, we optimize the validation of stored messages because the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
The number of proofs included in a blockchain transaction depends on the transaction cost limit set by the blockchain framework. Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added in a transaction payload $P_{j}$ until the limit is reached. Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording the accountable data to the blockchain.
@@
A service interruption or a disruption in the behavior of the autonomous agent will lead to the aggregation of the remaining hash values into a single transaction, even if there is capacity for additional values. This practice guarantees the immutability of the last messages sent by a robot upon completing a task or achieving a goal.
3.1.2 Smart Contract
Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm, effectively identifying the signer.
@@ 3.1.2 Smart Contract
Moreover, our SC includes a function to verify the existence of a message proof on the blockchain. This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored. This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
3.2 Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach is focused on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts relevant information. This functionality serves to enhance the effectiveness of explanations provided to non-expert users. A more comprehensive description of this component is provided in Section 4.
After processing the raw data stored by our accountability component, the ingestion process performs the chunking task to break down extensive text segments into more manageable units. To ensure comparability during relevancy searches at inference time, both this data and the user query are converted into numerical representations through embedding language models. Next, based on the embedding of the user query, its relevant text is identified in the document collection by a similarity search in the embedding space. The user-provided prompt is then extended with relevant information obtained in the search and added to the context. This enhanced prompt is sent to the LLM. As the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
The addition of new accountable data to our explainability component requires asynchronous updates of the accountable data and its embedding representation. These updates ensure that the model remains sensitive to changes, addressing limitations introduced by offline training.
4 Implementation
The solution presented in Section 3 can be implemented using different technologies. This section details the specific solutions chosen for developing our proposal.
4.1 Black Box Recorder
@@
Nodes constitute the fundamental components in both ROS and ROS 2 applications. These nodes are individual processes that handle specific tasks and communicate with each other through messages. Data transfer between nodes can be achieved using various methods, such as topics, services, and actions. Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that need feedback.
This work focuses on navigation capabilities. For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor to the ROS Navigation Stack. Nav2 equips the robot with the capacity to execute complex navigation tasks.
Nav2’s core functionalities include planning, control, localization, visualization, and additional features. Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and the specification of complex robot behaviors [84]. Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the BT over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [85].
Within ROS 2, nodes communicate by publishing data on topics, which can be consumed by other nodes. The data itself can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
Table 1: Topics name included in the Rosbag file their description.
@@
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain. Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data. Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work. In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability. The high-level Algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@
1:topicsSet = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
2:$rosbag$
3:Boolean $recording\leftarrow$ false
4:String $initialNonce\leftarrow$ sha256($randomBytes$)
5:// Recording process attending to a service call.
6:$recording\leftarrow serviceCallValue$
7:if $recording$ = $true$ then
8:     String $previousHash\leftarrow$$initialNonce$
9:     for $message$ in $topicsSet$ do
10:         if $messageNumber$ $topicRate$ == 0  then
11:              $chainedHash\leftarrow$ sha256($message$,$previousHash$)
12:              $previousHash\leftarrow$ $chainedHash$
13:         end if
14:         // Building chain of hashes.
15:         if $payloadSize$ < BLOCKS_LIMIT then
16:              $payload\leftarrow$ $payload\cup chainedHash$
17:         else
18:              // Load Solitidy contract.
19:              loadBcContract($contractAddress$, $contractAbi$)
20:              // Saving chain of hashes into blockchain.
21:              $transaction\leftarrow$buildTransaction($payload$)
22:              $signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
23:              sendTransaction($signed\_transaction$)
24:              for $hash$ in $payload$ do
@@
26:              end for
27:         end if
28:     end for
29:end if
30:return $rosbag$
Algorithm 2 Proof verification
1:$hashValue$
2:$BlockNumber$
3:// Load Solitidy contract.
4:loadBcContract($contractAddress$, $contractAbi$)
5:$transaction\leftarrow$buildTransaction($hashValue$)
6:$signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
7:sendTransaction($signed\_transaction$)
8:$blockNumber\leftarrow$readBC($hashValue$)
9:if $blockNumber$<>0 then
10:     $msg\leftarrow$ “The hash value is stored in $blockNumber$."
11:else
12:     $msg\leftarrow$ “The hash value is not stored."
13:end if
14:return $msg$
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of data while preserving system performance. These intervals define how often an integrity proof is stored and can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process. To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing an SC using the Solidity language. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number where it is stored. Verification of transaction ownership is also conducted during the storage of these proofs.
Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to address challenges related to the real-time recording and storage of Rosbag files securely and transparently.
4.2 Log Curation Component
To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented. Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process. By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
The ROS 2 NavigateToPose action server commands the robot to navigate to a specific goal. While Nav2 usually prioritizes the shortest path, external factors may require adjustments. The /plan topic provides the sequence of poses planned to reach the goal. If unexpected circumstances cause a deviation from this path, the Euclidean distance between consecutive poses in the precomputed plan will increase. This change, together with sensor data from the /scan topic, can indicate potential obstacles causing re-planning. These reasonings have been successfully checked in previous authors’ work [86] to build explanations for autonomous agents based on the use of ROS 2 topics.
Additionally, the present approach uses the information provided by Nav2 behavior trees due to their potential in designing, visualizing, and understanding the behavior of autonomous robots. Other information such as the present position of the robot and its velocity is also processed through this component. A general overview of its behavior is described through the Algorithm 3. As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without the need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of how raw ROS messages are transformed through this component is included in Figure 2.
Algorithm 3 Log Curation process
1:$rosbag$, $btNodeDesc$
2:$curationLog$
3:for $message$ in $rosbag$ do
4:     $msgTopic\leftarrow$ getMsgTopic($message$)
5:     $logMsg\leftarrow message.data$
6:     if $msgTopic$ == ’/plan’  then
7:         $logMsg\leftarrow$ Log significant path changes and obstacles.
8:     else if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
9:         $logMsg\leftarrow$ Track navigation status and log outcomes.
10:     else if $msgTopic$ == ’/behavior_tree_log’  then
11:         $logMsg\leftarrow$ Log behavior tree node status.
12:     else if $msgTopic$ == ’/amcl_pose’  then
13:         $logMsg\leftarrow$ Record robot position and orientation.
14:     else if $msgTopic$ == ’/cmd_vel’  then
15:         $logMsg\leftarrow$ Log linear and angular velocities.
@@ 4.2 Log Curation Component
{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’){mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Processed Data 1701354686 Navigation to the goal number 1 has started. 1701354686 Navigation to the goal number 1 is in progress. 1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
@@
4.3 Explainability Component
The output of the previously described node serves as an external source in our implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.

### v4

@@ III System Architecture
Our approach comprises two main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through the SC specifications. The explainability component provides natural language explanations by implementing RAG, thereby enhancing the results obtained using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
@@
Figure 1 depicts the generic architecture of our proposal, highlighting the functionalities related to accountability in blue, and those related to explainability in green. Below, we provide a deeper description of the approaches designed and followed to develop and integrate the above-mentioned components.
@@ III-A Black Box Recorder
The first component focuses on recording accountable information that captures the essential messages generated during the navigation task of an autonomous agent. The selected data will guarantee the reproducibility of the mentioned task, including details such as the robot’s poses, employed map, odometry data, linear and angular velocity, and laser scan readings.
@@
The immutability of the previously described information is achieved using blockchain. Blockchain technology usually faces limitations in terms of the size of the data that can be written into a single transaction. Efficient storage of autonomous agent messages is challenging owing to these constraints. Our proposed solution involves selectively storing the hashes of messages at intervals, providing the chance to use different storage frequencies depending on the topic and its relevance. This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
@@
Our anti-tampering approach to this accountability proposal comprises two key components: a chain of integrity proofs coupled with an SC specification. This section delves into their design and explains how both align with the constraints of the autonomous agents.
@@
Key features such as consistency are guaranteed by blockchain’s consensus protocol, in which the content of each block is hashed and stored in the next one. In addition, availability is assured by the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
III-A1 Chain of Integrity Proofs
@@
To preserve the integrity of the messages without compromising system performance, our solution entails building a chain where each message hash is linked with the previous hash value. This approach ensures that any alteration of the selected messages results in a different final hash, easing tampering detection. The verification process is also simplified because only the last hash in the chain must be recomputed and compared with the last stored value. A single hash modification disrupts the entire chain from that point onward, simplifying the identification of tampering.
@@
We define a chain entry ($CE$) with a counter $i\in 1..m$ to be linked with the previous one by concatenating the prior digest $h_{i-1}$ with the current log message $Msg_{i}$. hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, which is commonly used for data integrity verification and digital signatures.
@@
The initial hash value $h_{0}$ of the chain is calculated by applying the selected hash function to a set of $l=32$ random bytes ($RB$), to generate a 256-bit value. Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that the integrity proofs for similar initial log messages from different records are not identical. Data privacy is further prioritized by storing only the hashes in the blockchain, rather than the content of the messages.
@@
By chaining every hash with the previous one, we optimize the validation of the stored messages because the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
The number of proofs included in a blockchain transaction depends on the transaction cost limit set by the blockchain framework. Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added to a transaction payload $P_{j}$ until the limit is reached. Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording accountable data to the blockchain.
@@
A service interruption or a disruption in the behavior of the autonomous agent will lead to the aggregation of the remaining hash values into a single transaction, even if there is capacity for additional values. This practice guarantees the immutability of the last messages sent by the robot upon completing a task or achieving a goal.
III-A2 Smart Contract
Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm, effectively identifying the signer.
@@ III-A2 Smart Contract
Moreover, our SC includes a function to verify the existence of a message proof on the blockchain. This allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it is stored. This functionality contributes to the transparency of the system, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
III-B Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach focuses on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information is processed by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, which serves as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts the relevant information. This functionality enhances the effectiveness of the explanations provided to non-expert users. A more comprehensive description of this component is provided in Section IV.
After processing the raw data stored by our accountability component, the ingestion process performs a chunking task to break down extensive text segments into more manageable units. To ensure comparability during relevancy searches at inference time, both the data and user query are converted into numerical representations by embedding language models. Next, based on the embedding of the user query, the relevant text is identified in the document collection using a similarity search in the embedding space. The user-provided prompt is then extended with the relevant information obtained in the search and added to the context. This enhanced prompt is sent to the LLM. Because the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
The addition of new accountable data to our explainability component requires asynchronous updates of accountable data and its embedding representation. These updates ensure that the model is sensitive to changes, thereby addressing the limitations introduced by offline training.
IV Implementation
The solution presented in Section III can be implemented using various technologies. This section details the specific solutions used to develop our proposal.
IV-A Black Box Recorder
@@
Nodes constitute fundamental components of both ROS and ROS 2 applications. These nodes are individual processes that handle specific tasks and communicate with one another through messages. Data transfer between nodes can be achieved using various methods, such as topics, services, and actions. Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that require feedback.
This study focuses on navigation capabilities. For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor of the ROS Navigation Stack. Nav2 equips the robot with the capacity to execute complex navigation tasks.
Nav2’s core functionalities include planning, control, localization, visualization, and additional features. Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and specifications of complex robot behaviors [78]. Different independent task servers, intended to compute a path, control effort, or recovery, communicate with the BT over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has previously been successfully evaluated, allowing causal information to answer questions [79].
In ROS 2, nodes communicate by publishing data on topics that can be consumed by other nodes. The data can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes. The navigation reproducibility of autonomous robots in ROS 2 involves storing essential data such as their poses, maps, odometry, velocity, and laser scan readings. To this end, we used Rosbag files, designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding the system behavior. The set of topics included in the generated Rosbag files generated are detailed in Table I. The selected topics capture the core aspects of the robot’s operation, including localization, environment perception, navigation status, decision-making logic, and motion control. This selection ensures that the recorded data provides a coherent and sufficient basis for interpreting and reproducing the robot’s navigation behavior.
TABLE I: Topics name included in the Rosbag file their description.
@@
Ethereum has been selected as the blockchain framework to ensure the immutability of generated Rosbag files. Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain. Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without requiring approval, which is crucial for maintaining the integrity of the recorded data. Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as the Proof of Work. Furthermore, Ethereum’s scalability, security and economic efficiency make it a secure option for ensuring data immutability. High-level Algorithms 1 and 2 depict the main steps in the recording and proof verification services.
@@
0:  $topicsSet$ = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
0:  $rosbag$
1:  $recording\leftarrow$ false
2:  $initialNonce\leftarrow$ sha256($randomBytes$)
3:  $recording\leftarrow serviceCallValue$
4:  $proof\leftarrow\{\}$
5:  if $recording=true$ then
6:     $previousHash\leftarrow initialNonce$
7:     for $message$ in $topicsSet$ do
8:        if $messageNumber$ mod $topicRate$ = 0 then
9:           $chainedHash\leftarrow$ sha256($message$, $previousHash$)
10:           $previousHash\leftarrow chainedHash$
11:        end if
12:        if $payloadSize$ BLOCKS_LIMIT then
13:           $payload\leftarrow payload\cup chainedHash$
14:        else
15:           loadBcContract($contractAddress$, $contractAbi$)
16:           $transaction\leftarrow$ buildTransaction($payload$)
17:           $signedTx\leftarrow$ signTransaction($transaction$, $privateKey$)
18:           sendTransaction($signedTx$)
19:           for $hash$ in $payload$ do
20:              saveBC($proof[hash]$ = BLOCK_NUMBER)
21:           end for
22:        end if
23:     end for
24:  end if
25:  return  rosbag
Algorithm 2 Proof Verification
0:  $hashValue$, $contractAddress$, $contractAbi$, $privateKey$
0:  $msg$
1:  loadBcContract($contractAddress$, $contractAbi$)
2:  $transaction\leftarrow$ buildTransaction($hashValue$)
3:  $signedTx\leftarrow$ signTransaction($transaction$, $privateKey$)
4:  sendTransaction($signedTx$)
5:  $blockNumber\leftarrow$ readBC($hashValue$)
6:  if $blockNumber\neq 0$ then
7:     $msg\leftarrow$ “The hash value is stored in block ” $+blockNumber$
8:  else
9:     $msg\leftarrow$ “The hash value is not stored.”
10:  end if
11:  return  $msg$
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and the costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of the data while preserving system performance. These intervals define how often an integrity proof is stored and can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time in the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. The detection of any alteration to the messages within this file is facilitated by a service that generates a hash chain from the file at intervals set for each topic during the recording process. Blockchain logic has been extended by developing a SC using Solidity language to verify the existence of any of these hashes in the blockchain. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number in which it is stored. The verification of transaction ownership is also conducted during the storage of these proofs.
Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to securely and transparently address challenges related to the real-time recording and storage of Rosbag files.
IV-B Log Curation Component
To filter the relevant information from the Rosbag file messages, a raw data-processing node has been implemented. Given the size, variety, and intricate nature of the messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. To address this, a dedicated component has been developed to refine the data by discarding non-essential content before generating natural language explanations. Furthermore, by analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
The ROS 2 NavigateToPose action server commands the robot to navigate toward a specific goal. Although Nav2 usually prioritizes the shortest path, external factors may require adjustments. The /plan topic provides the sequence of poses planned to reach the goal. If unexpected circumstances cause a deviation from this path, the Euclidean distance between consecutive poses in the precomputed plan will increase. This change, together with sensor data from the /scan topic, can indicate potential obstacles that cause re-planning. These reasonings have been successfully verified in previous authors’ work [80] to build explanations for autonomous agents based on the use of ROS 2 topics.
Additionally, the present approach uses the information provided by Nav2 behavior trees because their potential for designing, visualizing, and understanding the behavior of autonomous robots. Other information, such as the present position of the robot and its velocity is also processed through this component. A general overview of its behavior is presented in Algorithm 3. As a result, through this component, ROS 2 developers are relieved from calling the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without the need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of the transformation of ROS messages through this component is shown in Figure 2.
Algorithm 3 Log Curation Process
0:  $rosbag$, $btNodeDesc$
0:  $curationLog$
1:  for $message$ in $rosbag$ do
2:     $msgTopic\leftarrow\text{getMsgTopic}(message)$
3:     $logMsg\leftarrow message.data$
@@ IV-B Log Curation Component
5:        $logMsg\leftarrow$ Track navigation status
[…]
7:        $logMsg\leftarrow$ Record log outcomes
[…]
9:        $logMsg\leftarrow$ Log path changes and obstacles
[…]
11:        $logMsg\leftarrow$ Log behavior tree node status
[…]
13:        $logMsg\leftarrow$ Record robot position and orientation
[…]
15:        $logMsg\leftarrow$ Log linear and angular velocities
@@
IV-C Explainability Component
The output of the previously described node serves as an external source in the implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact conversationally with documents while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.